OpenStudy (anonymous):

I need help with exponential regression...I have E(x)=164.96(1.03)^x for equation for high school graduates...need to find the first year of the decade in which number of graduates will reach 5 million

OpenStudy (anonymous):

E a function of x, E is what and x is what? im guessing E is graduates and x is year?

OpenStudy (anonymous):

yes I used cal to get a b and r

OpenStudy (anonymous):

plug 5mill into E(x) and solve for x

OpenStudy (anonymous):

yes got that but do I ln or log...can't remember

OpenStudy (anonymous):

years are in decades and graduates are in 1000s so would I divide 5 mil by 100

OpenStudy (anonymous):

do log base 1.03 there is probably a key on your calculator

OpenStudy (anonymous):

is that ln

OpenStudy (anonymous):

actually im sorry you are right take the natural log of both sides. and the exponent of x will come out as a multiple of ln(1.03) so \[\ln(1.03^x) = xln(1.03)\] ln(1.03) is just a constant so you can divide by it on both sides

OpenStudy (anonymous):

thank you for the confirmation

OpenStudy (anonymous):

yeahp

OpenStudy (anonymous):

one more thing I have year 2050 from the linear function and year 2070 from the exponential...which one seems more reasonable?

OpenStudy (anonymous):

based on the values of correlation factor r the linear function would be but if you're looking at 5 million wouldn't it take less years to get there Exponentially versus linearly

OpenStudy (anonymous):

im not sure from what was given i only know of E(x) and x,

OpenStudy (anonymous):

L(x)=33.79x-135.77

OpenStudy (anonymous):

used 5000 instead of 5 mil bc grads in 1000s is that right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah to which question, lol?

OpenStudy (anonymous):

lol just agreeing with the thing you said about linear versus exponential growth it would have to be 2050. but i would look into that more as to how to get to that answer

OpenStudy (anonymous):

ok thanks!